# Robust direction-dependent gain-calibration of beam-modelling errors far from the target field

### v1

@@ 3.6 Calibration
A telescope without missing elements is used to create model coherencies for calibration purposes. Because we have performed the forward prediction with unity gains for all except the switched-off elements, and because we do not assume band-pass errors, we only consider a DD-gain calibration step for source subtraction. If both direction-independent (DI) and DD-gain calibration are performed, errors in the DI gains could be absorbed in the DD-gains, such that the results become less interpretable. Calibration is performed in two steps, because the station beams change faster as a function of time and frequency for sources far from the target direction. Because of this, Cas A and Cyg A should be calibrated for and subtracted with a smaller time interval per solution than used for the target field [^13] In the LOFAR NCP field, these two steps are performed simultaneously with SAGECAL-CO. However, setting two different calibration intervals is not possible with the directionsolve method of DDECAL. (Gan et al., 2023; Munshi et al., 2024). The first round of calibration is done at time interval of 2.5 minutes per solution and in three directions: Cas A, Cyg A, and the target field for which we assume a single direction. Cas A and Cyg A are removed with the gain-calibration solutions for their respective directions applied to their respective sky-models. The residuals from this calibration step are used as an input for the second round of calibration using a solution interval of 20 minutes, in which the target field is divided into 24 gain-calibration directions. In the LOFAR-EoR key science project, DD-gain calibration is performed on a set of longer baselines ($250-5000\lambda$), to avoid absorption of the large-scale 21-cm signal in the calibration solutions (Patil et al., 2016; Mevius et al., 2022). The analysis is done on shorter baselines ($50-250\lambda$). Therefore, we also calibrate on these longer baselines only, but analyse all baselines between $50~{}\lambda$ and $5000~{}\lambda$. We omit three stations, that are a part of fewer than three baselines in this range. An overview of the calibration settings is listed in Table 2.
@@ 4.1 Gain Solutions
Because the Earth rotates during an observation, the spatial beam pattern shown in the last column of Figure 3 moves across the sky and changes shape (or, in the case of the NCP, rotates across the sky). Sources that are not in the station-beam main lobe pass through multiple nulls and sidelobes. Additionally, the position and width of the station-beam sidelobes depend on the spacing between antennas in units of wavelength. This causes the beam lobes to be narrower at higher frequencies, and wider at lower frequencies. These two effects combined define the beam as a four-dimensional structure with two spatial, one spectral and one temporal dimension. For a given point in the sky, this reduces to a two-dimensional dynamic spectrum for the gain.
Figure 4 shows the dynamic spectrum of the station beam from Figure 3 for the direction of Cas A. The top panel of Figure 4 shows the calibration beam (without switched-off elements), and the middle panel shows the true beam (with switched-off elements). When calibrating, the goal is to approximate the true beam with the product of the gain and calibration beam. The gain that would perfectly do this for a single point-source is equal to the ratio between the true beam and the calibration beam. This ideal gain is shown in the bottom panel of the figure.
@@
With a noisy 12 hr simulation in hand, we run DDECAL. The errors in the gain-calibration solutions in the direction of Cas A are shown in Figure 5. In all panels, the single point-source models are used for Cas A and Cyg A, but the spectral regularisation settings vary between panels. The gain error is computed by dividing the amplitude of the calculated gain by the amplitude of the ideal gain (i.e. the ratio of the true and calibration beams shown in the bottom panel of Figure 4). The top panel of Figure 5 shows the result if no spectral regularisation is used during calibration. This calibration method is not useful in practice, because it leads to overfitting of the data. This unregularised solution is therefore only shown for illustration purposes. The discretisation of the gains, discussed above, is clearly visible as steps in the gains every 2.5 minutes. Additionally, the gains are underestimated along the null in the calibration beam, as is evident from the orange regions in the same areas where the null was identified in Figure 4. Because the ideal gain approaches infinity here, this underestimation is expected. However, because the null in the calibration beam is narrow, this affects few channels, that could be flagged later.
@@ 4.2 Residual visibilities
We now turn our attention to the residual visibilities that are found after subtracting all sources with their respective DD-calibrated gains. Because there are no differences between the simulated and calibration sky models, a perfect calibration would result in the residual visibilities being completely noise-like [^14] Due to this being a non-linear model, a small bias is expected, but this is negligible compared to the error level.. The accuracy of the calibration can therefore be tested by comparing the residual visibilities to noise. In Figures 6 and 7, the dynamic spectra of the residual visibilities on single baselines are shown. One of the stations that comprise the baseline in Figure 6 is the same station as used for Figures 3, 4 and 5, such that the nulls seen in Figures 4 and 5 can be associated with effects seen in Figure 6. The left-hand columns of Figures 6 and 7 show the residual visibilities after subtracting the gain-calibrated sky model with a 4 MHz kernel and the standard method, and the right-hand columns show the same, but with the robust gain-calibration method. The top panels show visibility amplitudes, and the bottom panels show visibility phases.
@@
Figure 6: Residual visibilities for the baseline between CS001HBA0 and CS003HBA1 on 28-06-2014 (N3) after all sources have been subtracted with solutions created with the 4 MHz kernel. The times and frequencies match Figures 4 and 5. Left: Standard DDECAL algorithm, right: the robust method. Top: amplitudes, bottom: phases. Ideally, these residuals should be noise-like.
@@
Because perfect calibration would result in only Gaussian noise, the effectiveness of the calibration can be judged by deviations from the thermal noise level. Therefore, we compare the standard deviation of a complete 12 hr simulation before and after outlier excision to the thermal noise level of $\sigma_{n}=SEFD/\sqrt{\Delta\nu\Delta t}=3.01$ Jy. The standard deviations of the simulations of N1 (that starts on 27-12-2013, see Table 1) with the single-point and multi-point models for Cas A and Cyg A are shown in Figure 8. The figure shows the ratio between the standard deviations and noise before and after outlier flagging in the top and middle row respectively. The bottom row shows the percentage of data that was excised by AOFLAGGER. We consider the latter statistic, because using fewer flags leads to a lower loss of data, and therefore, a higher sensitivity. This loss of data can be significant, e.g. in LOFAR-EoR science, through various methods of flagging (see Mertens et al. 2025), up to 19% of data are flagged to excise Cas A and Cyg A (Florent Mertens, private communications). Therefore, attaining the same standard deviation with fewer flags is beneficial.
@@
The different panel columns show various kernel sizes for the calibration. The first and second columns show the 1 MHz and 4 MHz kernels, respectively. Similar to what was found in Section 4.1, the 4 MHz kernel performs significantly worse than the 1 MHz kernel. This is somewhat surprising, since Gan et al. (2023) showed that the 4 MHz kernel performed better on LOFAR-EoR data. We suspect that there is a difference in which kernel width performs better on the main lobe versus off-axis sources. Because the target field does not suffer from beam modelling errors to the same degree as the off-axis sources, the 4 MHz kernel may outperform the 1 MHz kernel for the target field sources, while this is the other way around for the far-field sources. Because the target field contributes more power than the off-axis sources, owing to lower beam attenuation, it dominates the residuals in real data, where calibration and subtraction are less ideal than in these simulations. To test this hypothesis, we also use a combined kernel on the simulation: a 1 MHz kernel is used for subtracting Cas A and Cyg A, whereas a 4 MHz kernel is used for the target field. The results for this dynamic calibration kernel are shown in the final column of Figure 8 and are very close to the results for the 1 MHz kernel. This suggests that the outliers found in the data primarily correspond to errors made in the directions of Cas A and Cyg A, such that these dominate the statistics shown in the figure.
@@
Figure 8 shows two sets of baselines: those between 250 $\lambda$ and 5000 $\lambda$ (i.e. the calibration baselines), and those between 50 $\lambda$ and 250 $\lambda$ (i.e. the EoR analysis baselines). Comparing the calibration and analysis baselines provides information about the level of overfitting, ideally the two sets should be as close as possible. However, there is more residual power on the analysis baselines. Furthermore, the analysis baselines in the standard method require a higher flagging percentage than those in the robust method.
@@
The horizontal axis coordinates within a panel of Figure 8 denote two different models for the off-axis sources, describing them either with a single or multiple point sources. Because the LSTs are identical between the simulations, only the off-axis contribution to the visibilities differs, whereas that of the target field stays the same. Clearly, the single point-source model suffers more from outliers than the multi-point source model, as is evident by the reduction in standard deviation during flagging and the number of flags. This may be due to two things: firstly, the multi-point source model has a bigger spatial footprint, and secondly, each individual component of the multi-point source model contains less power than the single component of the single-point model. To differentiate between these two, we also performed a test with a strongly attenuated single-point model (with the flux reduced to 5 kJy for Cyg A and 4.2 kJy for Cas A). This attenuated model suffers from large outliers in the same way as the single point-source model does (see the complete results in Figure 10), such that we can attribute the better calibration behaviour of the multi-point source model to its bigger spatial footprint. As a result of the extended spatial footprint, the source is never fully inside a null, such that its flux is not as abruptly attenuated and the spectral smoothness assumption is not violated as strongly.
@@ 4.3 Imaging
To this end, the residual data is continuum imaged at a resolution of 0.3 arcmin with a field of $10^{\circ}\times 10^{\circ}$ and uniform weighting using WSCLEAN (Offringa et al., 2014). All baselines between 50 $\lambda$ and 5000 $\lambda$ are used in the images. To give a realistic comparison, flags are applied to the data before imaging, such that the outliers are excised. As a result, none of the images contain very bright sources anymore, and the dirty images can be analysed directly, without a need for deconvolution. Subsequently, the images are Fourier transformed to determine their angular power spectrum (i.e. the power at various angular scales). If there is residual source power in the images due to gain-calibration errors, this will show as additional power above the thermal noise level. Therefore, better DD-gain calibration generally leads to less power in image space, although sources may leave more residual power on some spatial scales than others. Figure 9 shows the ratio between the power spectrum of the images created using the robust and standard regularisation methods as a function of spatial scale. Points below unity mean that the robust regularisation method performs better. The left panel shows the results with the point-source model (for all nights), and the right panel shows the results with the multi-point model. The different colors indicate different smoothing kernels.
@@
Similar to what was found in visibility space, the robustly regularised DD-gain calibration method performs similarly to or better than the standard method on throughout the spectrum. In simulation N3 with the 4 MHz kernel, robust regularisation performs especially well compared to standard regularisation. We expect this to be because the off-axis sources have a higher elevation at the LST of this simulation. Because of this, they are less attenuated by the antenna beam for this simulation and, therefore, have the highest impact. A similar effect is found for the standard deviation of this simulation in Appendix A. This suggests that the robust regularisation method is able to mitigate the impact of strong sources in the sidelobes better on all scales in image space, even when strong spectral regularisation is used.
@@
Overall, however, the ratio remains near unity for both model types, except at the lowest $\ell$-modes, especially those related to scales $\gtrsim$0.06 arcmin-1. These largest scales in the image correspond to the baselines with a projected length of $\leq 250~{}\lambda$, at which the robust method performs significantly better than the standard method. These scales are important for many science cases, such as 21-cm signal observations of the Epoch of Reionization and Cosmic Dawn, and studies of the diffuse polarised foreground of the Milky Way. We attribute the improvement on the short baselines to the distinction between calibration and analysis baselines. This implies that the robust spectral regulariser results in gains that extrapolate better to shorter baselines than the standard regulariser, which may occur because the gains are less biased by strong spectral variations in the beam-sidelobes.
@@ 5 Discussion and conclusions
In conclusion, beam modelling for phased-array stations such as those of LOFAR, NenuFAR, MWA (where we consider a tile to be a station) and the upcoming SKA remains a challenging endeavour due to changes in the beam shapes as a function of pointing, time (due to rotation of the Earth), and observing frequency, as well as additional temporal changes due to the behaviour of the telescope due to degrading hardware and weather. The upcoming SKA-Low array will have a higher sensitivity than the current generation of instruments, and likely also a more complex antenna beam. Therefore, beam effects such as those present for LOFAR-HBA stations will similarly affect SKA. Furthermore, SKA-Low’s higher sensitivity may cause far-field sources currently below the noise to appear on the data. Therefore, applying DD-gain calibration methods that are robust against beam-modelling errors in the distant sidelobes and nulls are necessary to use radio telescopes to their full capabilities.

### v2

@@ 3.6 Calibration
A telescope without missing elements is used to create model coherencies for calibration purposes. Because we have performed the forward prediction with unity gains for all except the switched-off elements, and because we do not assume band-pass errors, we only consider a DD-gain calibration step for source subtraction. If both direction-independent (DI) and DD-gain calibration are performed, errors in the DI gains could be absorbed in the DD-gains, such that the results become less interpretable. Calibration is performed in two steps, because the station beams change faster as a function of time and frequency for sources far from the target direction. Because of this, Cas A and Cyg A should be calibrated for and subtracted with a smaller time interval per solution than used for the target field [^14] In the LOFAR NCP field, these two steps are performed simultaneously with SAGECAL-CO. However, setting two different solution intervals is not possible with the directionsolve method of DDECAL. (Gan et al., 2023; Munshi et al., 2024). The first round of calibration is done at time interval of 2.5 minutes per solution and in three directions: Cas A, Cyg A, and the target field for which we assume a single direction. Cas A and Cyg A are removed with the gain-calibration solutions for their respective directions applied to their respective sky-models. The residuals from this calibration step are used as an input for the second round of calibration using a solution interval of 20 minutes, in which the target field is divided into 24 gain-calibration directions. In the LOFAR-EoR key science project, DD-gain calibration is performed on a set of longer baselines ($250-5000\lambda$), to avoid absorption of the large-scale 21-cm signal in the calibration solutions (Patil et al., 2016; Mevius et al., 2021). The analysis is done on shorter baselines ($50-250\lambda$). Therefore, we also calibrate on these longer baselines only, but analyse all baselines between $50~{}\lambda$ and $5000~{}\lambda$. We omit three stations, that are a part of fewer than three baselines in this range. An overview of the calibration settings is listed in Table 2.
@@ 4.1 Gain Solutions
Because the Earth rotates during an observation, the spatial beam pattern shown in the last column of Figure 3 moves across the sky and changes shape (or, in the case of the NCP, rotates across the sky). Sources that are not in the station-beam main lobe pass through multiple nulls and sidelobes. Additionally, the position and width of the station-beam sidelobes depend on the spacing between antennas in units of wavelength. This causes the beam lobes to be narrower at higher frequencies, and wider at lower frequencies. These two effects combined define the beam as a four-dimensional structure with two spatial, one spectral and one temporal dimension. For a given point in the sky, this reduces to a two-dimensional dynamic spectrum for the gain $|\bm{\mathsf{J}}_{p,k}(\nu,t)|$.
Figure 5 shows the dynamic spectrum of the station power beam [^15] As is customary in literature about beams, we use a ‘power beam’ in the plots. The power beam is $|\bm{\mathsf{A}}_{p,k}(\nu,t)|^{2}$. from Figure 3 for the direction of Cas A. The top panel of Figure 5 shows the calibration beam (without switched-off elements), and the middle panel shows the true beam (with switched-off elements). When calibrating, the goal is to approximate the true beam with the product of the gain and calibration beam. The gain that would perfectly do this for a single point-source is equal to the ratio between the true beam and the calibration beam. This ideal gain is shown in the bottom panel of the figure.
@@
With a noisy 12 hr simulation in hand, we run DDECAL. The errors in the gain-calibration solutions in the direction of Cas A are shown in Figure 6. In all panels, the single point-source models are used for Cas A and Cyg A, but the spectral regularisation settings vary between panels. The gain error is computed by dividing the amplitude of the calculated gain by the amplitude of the ideal gain (i.e. the ratio of the true and calibration beams shown in the bottom panel of Figure 5). The top panel of Figure 6 shows the result if no spectral regularisation is used during calibration, and generally follows the expected gain behaviour shown in the bottom panel of Figure 5. The unregularised solution is not useful in practice due to overfitting of the data, as is visible as the rapid variations in the lower right corner of the plot. We include the unregularised solution here purely as a reference plot, to illustrate the effect that regularisation has. The discretisation of the gains, discussed above, is clearly visible as steps in the gains every 2.5 minutes. Additionally, the gains are underestimated along the null in the calibration beam, as is evident from the orange regions in the same areas where the null was identified in Figure 5. Because the ideal gain approaches infinity here, this underestimation is expected. However, because the null in the calibration beam is narrow, this affects few channels, that could be flagged later.
@@ 4.2 Residual visibilities
We now turn our attention to the residual visibilities that are found after subtracting all sources with their respective DD-calibrated gains. Because there are no differences between the simulated and calibration sky models, a perfect calibration would result in the residual visibilities being completely noise-like [^16] Due to this being a non-linear model, a small bias is expected, but this is negligible compared to the error level.. The accuracy of the calibration can therefore be tested by comparing the residual visibilities to noise. In Figures 7 and 8, the dynamic spectra of the residual visibilities on single baselines are shown. One of the stations that comprise the baseline in Figure 7 is the same station as used for Figures 3, 5 and 6, such that the nulls seen in Figures 5 and 6 can be associated with effects seen in Figure 7. The left-hand columns of Figures 7 and 8 show the residual visibilities after subtracting the gain-calibrated sky model with a 4 MHz kernel and the standard method, and the right-hand columns show the same, but with the robust gain-calibration method. The top panels show visibility amplitudes, and the bottom panels show visibility phases.
@@
Figure 7: Residual visibilities for the baseline between CS001HBA0 and CS003HBA1 on 28-06-2014 (N3) after all sources have been subtracted with solutions created with the 4 MHz kernel. The times and frequencies match Figures 5 and 6. Left: Standard DDECAL algorithm, right: the robust method. Top: amplitudes, bottom: phases. Ideally, these residuals should be noise-like.
@@
Because perfect calibration would result in only Gaussian noise, the effectiveness of the calibration can be judged by deviations from the thermal noise level. Therefore, we compare the standard deviations of the visibilities of a complete 12 hr simulation before and after outlier excision to the thermal noise level of $\sigma_{n}=SEFD/\sqrt{\Delta\nu\Delta t}=3.01$ Jy. The standard deviations of the simulations of N1 (that starts on 27-12-2013, see Table 1) with the single-point and multi-point models for Cas A and Cyg A are shown in Figure 9. The figure shows the ratio between the standard deviations and noise before and after outlier flagging in the top and middle row respectively. The bottom row shows the percentage of data that was excised by AOFLAGGER. We consider the latter statistic, because using fewer flags leads to a lower loss of data, and therefore, a higher sensitivity. This loss of data can be significant, e.g. in LOFAR-EoR science, through various methods of flagging (see Mertens et al. 2025), up to 19% of data are flagged to excise Cas A and Cyg A (Florent Mertens, private communications). Therefore, attaining the same standard deviation with fewer flags is beneficial.
@@
The different panel columns show various kernel sizes for the calibration. The first and second columns show the 1 MHz and 4 MHz kernels, respectively. Similar to what was found in Section 4.1, the 4 MHz kernel performs significantly worse than the 1 MHz kernel. This is somewhat surprising, since Gan et al. (2023) showed that the 4 MHz kernel performed better on LOFAR-EoR data. We suspect that there is a difference in which kernel width performs better on the main lobe versus off-axis sources. Because the target field does not suffer from beam modelling errors to the same degree as the off-axis sources, the 4 MHz kernel may outperform the 1 MHz kernel for the target field sources, while this is the other way around for the far-field sources. A similar spectral smoothness difference between the main lobe and off-axis sources is also observed in NenuFAR observations (Munshi et al., 2025). For both instruments, the target field contributes more power than the off-axis sources, owing to lower beam attenuation. Because of this, the target field dominates the residuals in real data, where calibration and subtraction are less ideal than in these simulations. In our simulations, the only errors that are introduced are beam modelling errors and thermal noise. Because the beam modelling errors are spectrally smooth within the main lobe of the primary beam, we are able to remove the target field sources to a much deeper level, revealing errors in the directions of Cas A and Cyg A.
To test this hypothesis, we also use a combined kernel on the simulation: a 1 MHz kernel is used for subtracting Cas A and Cyg A, whereas a 4 MHz kernel is used for the target field. The results for this dynamic calibration kernel are shown in the final column of Figure 9 and are very close to the results for the 1 MHz kernel. This suggests that the outliers found in the data primarily correspond to errors made in the directions of Cas A and Cyg A, such that these dominate the statistics shown in Figure 9.
@@
Figure 9 shows two sets of baselines: those between 250 $\lambda$ and 5000 $\lambda$ (i.e. the calibration baselines), and those between 50 $\lambda$ and 250 $\lambda$ (i.e. the EoR analysis baselines). Comparing the calibration and analysis baselines provides information about the level of overfitting, ideally the two sets should be as close as possible. However, there is more residual power on the analysis baselines. This is expected for a baseline cut, and corresponds to earlier findings (Mevius et al., 2021). Furthermore, the analysis baselines in the standard method require a higher flagging percentage than those in the robust method.
@@
The horizontal axis coordinates within a panel of Figure 9 denote two different models for the off-axis sources, describing them either with a single or multiple point sources. Because the LSTs are identical between the simulations, only the off-axis contribution to the visibilities differs, whereas that of the target field stays the same. Clearly, the single point-source model suffers more from outliers than the multi-point source model, as is evident by the reduction in standard deviation during flagging and the number of flags. This may be due to two things: firstly, the multi-point source model has a bigger spatial footprint, and secondly, each individual component of the multi-point source model contains less power than the single component of the single-point model. To differentiate between these two, we also performed a test with a strongly attenuated single-point model (with the flux reduced by a factor four). This attenuated model suffers from large outliers in the same way as the single point-source model does (see the complete results in Figure 11), such that we can attribute the better calibration behaviour of the multi-point source model to its bigger spatial footprint. As a result of the extended spatial footprint, the source is never fully inside a null, such that its flux is not as abruptly attenuated and the spectral smoothness assumption is not violated as strongly.
@@ 4.3 Imaging
[…]
$P(\ell)=\left<\left|\mathcal{F}\left\{I(x,y)\right\}\right|^{2}\right>_{\ell}.$ (17)
Here, ${}\mathcal{F}\left\{\cdot\right\}$ denotes the spatial Fourier transform, and $I(x,y)$ denotes the image intensity at pixel $(x,y)$. The operator $_{\ell}$ represents an azimuthal average over Fourier modes with constant angular frequency $\ell$. If $\ell_{x}$ and $\ell_{y}$ are the Fourier conjugates of the image axes, $\ell=\sqrt{\ell_{x}^{2}+\ell_{y}^{2}}$. We use the cosmological convention of using the inverse of the spatial scale as the Fourier conjugate.
@@
If there is residual source power in the images due to gain-calibration errors, this will show as additional power above the thermal noise level. Therefore, better DD-gain calibration generally leads to less power in image space, although sources may leave more residual power on some spatial scales than others. Figure 10 shows the ratio between the power spectrum of the images created using the robust and standard regularisation methods as a function of spatial scale, i.e. $P_{\mathrm{robust}}(\ell)/P_{\mathrm{standard}}(\ell)$, where the subscripts ‘robust’ and ‘standard’ denote the calibration methods. Points below unity generally mean that the robust regularisation method performs better, although the ratio also contains noise. The left panel shows the results with the point-source model (for all nights), and the right panel shows the results with the multi-point model. The different colours indicate different smoothing kernels.
@@
Similar to what was found in visibility space, the robustly regularised DD-gain calibration method performs similarly to or better than the standard method throughout the spectrum, with only a few points exceeding unity, which we attribute to noise. In simulation N3 with the 4 MHz kernel, robust regularisation performs especially well compared to standard regularisation. We expect this to be because the off-axis sources have a higher elevation at the LST range of this simulation. Because of this, they are less attenuated by the antenna beam for this simulation and, therefore, have the highest impact. A similar effect is found for the standard deviation of this simulation in Appendix A. This suggests that the robust regularisation method is able to mitigate the impact of strong sources in the sidelobes better on all scales in image space, even when strong spectral regularisation is used.
@@
Overall, however, the ratio remains near unity for both model types, except at the lowest $\ell$-modes, especially those related to scales $\lesssim$0.06 arcmin-1. These largest scales in the image correspond to the baselines with a projected length of $\leq 250~{}\lambda$, at which the robust method performs significantly better than the standard method. These scales are important for many science cases, such as 21-cm signal observations of the Epoch of Reionization and Cosmic Dawn, and studies of the diffuse polarised foreground of the Milky Way. We attribute the improvement on the short baselines to the distinction between calibration and analysis baselines. This implies that the robust spectral regulariser results in gains that extrapolate better to shorter baselines than the standard regulariser171717due to the baseline cut, short baseline-gains are constructed from the gain-solutions found on longer baselines., which may occur because the gains are less biased by strong spectral variations in the beam-sidelobes.
@@ 5 Discussion and conclusions
The robust regulariser has not been tested in the presence of more general cases of extended emission. This is because current DDECAL evaluates the beam response of extended emission at discrete points, effectively removing the difference between point sources and extended sources in the context of beam modelling errors. Furthermore, beam modelling errors are expected to have a lower impact on extended emission, since extended emission does not fully fall into deep beam nulls. Therefore, the difference between the standard and robust regulariser is expected to be smaller for extended emission. Using the difference between the single-point and multi-point off-axis source models as a first-order probe for how the robust regulariser performs for extended emission, we expect extended emission to stabilise the solutions further.
Diffuse extended emission in the sidelobes of the beam has been shown to pose a challenge for 21-cm cosmology with the drift-scan instrument HERA (N. Charles et al., 2023), necessitating a spatial filtering technique to remove this emission prior to calibration. This is not expected to be needed for LOFAR, because LOFAR has more long baselines, allowing it to have a higher baseline cut, that drastically reduces the impact of diffuse emission during gain-calibration. The HERA baseline cut is at 40 m, whereas that of LOFAR is at $250~{}\lambda\approx 500~{}$m in the relevant frequency range. Furthermore, Höfer et al. (2025) have shown that the combination of beam errors and diffuse emission does not lead to a bias in the 21-cm power spectrum for LOFAR during direction-independent calibration with a lower baseline cut.
In conclusion, beam modelling for phased-array stations such as those of LOFAR, NenuFAR, MWA (where we consider a tile to be a station) and the upcoming SKA-Low remains a challenging endeavour due to changes in the beam shapes as a function of pointing, time (due to rotation of the Earth), and observing frequency, as well as additional temporal changes due to the behaviour of the telescope due to degrading hardware and weather. The SKA-Low array, for example, will have a higher sensitivity than the current generation of instruments, and likely also have a more complex antenna beam. Therefore, beam effects such as those present for LOFAR-HBA stations will similarly affect SKA-Low. Furthermore, SKA-Low’s higher sensitivity may cause far-field sources currently below the noise to appear in the data. Therefore, applying DD-gain calibration methods that are robust against often unavoidable beam-modelling errors in the distant sidelobes and nulls are necessary to use phased-array telescopes to their full capabilities.